OpenStudy (anonymous):

What equation has a graph that is perpendicular to the graph of -x+6y=-12. Can you please show me how this is done

OpenStudy (123ab456c):

hi i can be a little help!:)

OpenStudy (mathmale):

1. Find the slope of -x+6y=-12. 2. Find the negative reciprocal of this line, to pin down the slope of the perpendicular line. 3. Write (in slope-intercept form) the equation y=mx+b. Choose any value you want for b. m, the slope, must be the negative reciprocal of the slope of -x+6y=-12.

OpenStudy (anonymous):

There are several answers. I am trying to help my son and am so confusedWhich equation has a graph that is perpendicular to the graph of −x + 6y = −12? A. x + 6y = −67 B. x − 6y = −52 C. 6x + y = −52 D. 6x − y = 52

OpenStudy (anonymous):

I cannot figure out how to get the answer. It's been a long time

OpenStudy (anonymous):

I am getting y=1/6x-2 so then would the negative reciprical be -6 OMG so confused

OpenStudy (mathmale):

Solve −x + 6y = −12 for y: First, add x to both sides to move the -x to the right side of the equation. You should get 6y=x-12. Your slope is correct: 1/6. Yes, and the neg recip is indeed (-6). y=mx + b. Use -6 for the slope m and choose any value you wish for b. That'll give you the answer you and your son wanted.

OpenStudy (anonymous):

I have absolutely no clue where they are getting the 52 or the 67

OpenStudy (anonymous):

we chose B, but that was incorrect

OpenStudy (mathmale):

Remember, the value of b is arbitrary. You need to choose the answer that is most likely. Answers A and B have the wrong slope. You have y =-6x + b. Move the 6x to the left side by adding 6x to both sides. 6x+y=b. b is arbitrary. Which answer choice is correct?

OpenStudy (anonymous):

so, the answer would be C?

OpenStudy (mathmale):

Yes, indeed. Nice work. Discuss our reasoning with your son. Wish I were there to help with that. Have to get off the 'Net now, but will be on again later tonight. Take care.

OpenStudy (anonymous):

Thanks! I was confusing myself with the numbers!!! Appreciate the help!

OpenStudy (mathmate):

For your information, when the given equation is in the form Ax+By+C=0 or Ax+By=C, A line perpendicular to the given can be found by 1. switch the coefficients of x and y 2. change the sign of ONE of the coefficients. Example: Find a line perpendicular to 4x-3y=8...........(1) A line perpendicular to (1) is 3x+4y=8...........(2) The constant term 8 does not change the slope. Sometimes you need to adjust the constant term to pass through a given point.

OpenStudy (anonymous):

Thanks!